OpenStudy (anonymous):

SOMEBODY HELP ME WITH THIS PLEASE!! What is the function of the equation (input/output) y = -5x - 3

OpenStudy (anonymous):

It is a linear equation

OpenStudy (anonymous):

I don't know how to solve and put in the table input and output

OpenStudy (anonymous):

x y 0 -3 y =-5(0)-3 choose a value for x, substitute it into the equation, y = -3 then solve for y 1 -8 y = -5(1)-3 y = -8 -1 2 y = -5(-1)-3 y = 2

OpenStudy (anonymous):

ok so how I put this in the table of function input/output

OpenStudy (anonymous):



OpenStudy (anonymous):

Thank you very much

OpenStudy (anonymous):

You're welcome